OpenStudy (wwwhhhaaattt?):

Will Fan and medal... DUE IN A COUPLE MINS 4^-3 x 4^6

OpenStudy (anonymous):

hold on please

OpenStudy (anonymous):

4^-3=(1/64) 4^6=4096 (1/64)*4096=64

OpenStudy (anonymous):

64 u can go on google and put that question in and get the answer by using the calcator